Which of the following options is better and why? Assume an opportunity cost of 12.5%

 Option A: Receive $10,000 today

 Option B: Receive $15,000 five years from today
Answer #1

In order to compute which of the two option is better, we need to apply time value of money concept to bring both cash flows at same point in time.

FV = PV * (1 + r)n

For option 2,

$15000 = PV * (1 + 12.5%)5

$15000 = PV * 1.8020

PV = $8,323.93 --> This is less than the option 1. Hence choose option 1 as that is more valuable.